# Feferman’s completeness theorem

### v1

@@ 5. Lower bound for Feferman’s completeness
Let $\mathcal{C}$ be the standard order of order type $\varepsilon_{0}$, whose elements are nested Cantor normal form terms. For $\alpha<\varepsilon_{0}$ we put $\mathsf{RFN}^{\alpha}(T)$ to be $\mathsf{RFN}^{(\mathcal{C},\alpha)}(T)$.
@@
contains $\mathsf{RFN}^{\mathcal{L}}(\mathsf{PA})$, for any $\mathsf{PA}$-verifiable computable well-ordering $\mathcal{L}$ of order type $\leq\omega^{n}$. However, the arithmetically sound theory $T$ has a $\Sigma_{2n}$-set of theorems and thus there is a true $\Pi_{2n}$ sentence $\varphi$ that is not provable in this theory. Hence $\varphi$ can not be proved by any $\mathsf{RFN}^{\mathcal{L}}(\mathsf{PA})$ of the considered form. ∎
@@ 6. Reflection iterated along elements of $\mathcal{O}$
A testament that Section 1.1 cannot be transferred to work with progressions along ordinal notation systems is that it gives for every true $\Pi_{1}$ sentence $\varphi$ a computable well-order $\mathcal{L}$ of order type $1$ such that $\mathsf{RFN}^{\mathcal{L}}(\mathsf{PA})\vdash\varphi$. Ordinal notation systems such as Kleene’s $\mathcal{O}$ have a unique notation for finite orderings and thus the analogous theorem for ordinal notation systems must fail. However, we can still obtain tight bounds for progressions along ordinal notation systems by using basic properties of computable linear orderings.
@@
Let $\varphi$ be a $\Pi_{2n+1}$ sentence and $\mathcal{L}$ be the linear ordering obtained in Corollary 16. Then $\mathsf{ACA}_{0}$ proves $\mathsf{WO}(\mathcal{L})$ if and only if $\varphi$ is true. In particular, the ordering $\mathcal{L}^{\prime}=\omega\cdot(1+\mathcal{L})+1$ is a computable ordering such that $\mathsf{ACA_{0}}$ proves $\mathcal{L}^{\prime}\cong\omega^{n+1}+1$ if and only if $\varphi$ is true. We can also assume that $\mathcal{L}^{\prime}$ comes with other special properties shared with notations in Kleene’s $\mathcal{O}$: It has a computable successor relation, a computable set of limit points and there are algorithms that compute first and last elements, predecessors, and fundamental sequences for limit points. For example, the ordering defined by
@@
Corollary 20 (Feferman’s original completeness theorem).
@@
A similar result was claimed in [fenstad1968], however Savitt found a mistake in this proof [savitt1969jens].
@@ 7. Discussion and Open Questions
Although we haven’t done this in the paper, it appears to us that it is easy to adopt the techniques from Section 1.1 to show sharpness of the upper bound from Theorem 19 on the order types for the iterations of reflection along elements of $\mathcal{O}$ necessary to attain the completeness of iterations of reflection for $\Pi_{2n+1}$-sentences.
In the paper we have been focusing on the iterations of full uniform reflection, however when we want to obtain some true $\Pi_{n}$-sentence in principle it is natural to expect that we need to iterate just partial uniform $\Pi_{n}$ reflection $\mathsf{RFN}_{\Pi_{n}}$ instead of full uniform reflection $\mathsf{RFN}$. For example, the existence of such iterations could be easily proved from combining Feferman’s completeness theorem for full uniform reflection together with Schmerl’s formula that allows to transform shorter iteration of stronger reflection to a longer iteration of weaker $\Pi_{n}$ reflection, while preserving the set of $\Pi_{n}$-consequences. It appears to us that the methods used in this paper can be adopted to show that for some natural number constant $C$ and every $n$ we can replace $\mathsf{RFN}$ with $\mathsf{RFN}_{\Pi_{2n+1+C}}$ in Section 1.1. However, it is not clear to us whether in Section 1.1 full uniform reflection $\mathsf{RFN}$ can be replaced with just $\mathsf{RFN}_{\Pi_{2n+1}}$, while keeping the bound $\omega^{n}+1$ on the order type.

### v2

@@ 5. Lower bound for Feferman’s completeness
Let $\mathcal{C}$ be the standard order of order-type $\varepsilon_{0}$, whose elements are nested Cantor normal form terms. For $\alpha<\varepsilon_{0}$ we put $\mathsf{RFN}^{\alpha}(T)$ to be $\mathsf{RFN}^{(\mathcal{C},\alpha)}(T)$.
@@
contains $\mathsf{RFN}^{\mathcal{L}}(\mathsf{PA})$, for any $\mathsf{PA}$-verifiable computable well-ordering $\mathcal{L}$ of order-type $\leq\omega^{n}$. However, the arithmetically sound theory $T$ has a $\Sigma_{2n}$-set of theorems and thus there is a true $\Pi_{2n}$ sentence $\varphi$ that is not provable in this theory. Hence $\varphi$ can not be proved by any $\mathsf{RFN}^{\mathcal{L}}(\mathsf{PA})$ of the considered form. ∎
@@ 6. Reflection iterated along elements of $\mathcal{O}$
A testament that Section 1.1 cannot be transferred to work with progressions along ordinal notation systems is that it gives for every true $\Pi_{1}$ sentence $\varphi$ a computable well-order $\mathcal{L}$ of order-type $1$ such that $\mathsf{RFN}^{\mathcal{L}}(\mathsf{PA})\vdash\varphi$. Ordinal notation systems such as Kleene’s $\mathcal{O}$ have a unique notation for finite orderings and thus the analogous theorem for ordinal notation systems must fail. However, we can still obtain tight bounds for progressions along ordinal notation systems by using basic properties of computable linear orderings.
@@
Let $\varphi$ be a $\Pi_{2n+1}$ sentence and $\mathcal{L}$ be the linear ordering obtained in Corollary 17. Then $\mathsf{ACA}_{0}$ proves $\mathsf{WO}(\mathcal{L})$ if and only if $\varphi$ is true. In particular, the ordering $\mathcal{L}^{\prime}=\omega\cdot(1+\mathcal{L})+1$ is a computable ordering such that $\mathsf{ACA_{0}}$ proves $\mathcal{L}^{\prime}\cong\omega^{n+1}+1$ if and only if $\varphi$ is true. We can also assume that $\mathcal{L}^{\prime}$ comes with other special properties shared with notations in Kleene’s $\mathcal{O}$: It has a computable successor relation, a computable set of limit points, and there are algorithms that compute the first and last elements, predecessors, and fundamental sequences for limit points. For example, the ordering defined by
@@
Corollary 21 (Feferman’s original completeness theorem with improved bounds).
@@
A similar result was claimed in [fenstad1968]; however, Savitt found a mistake in this proof [savitt1969jens].
@@ 7. Discussion and Open Questions
We suspect that it should be easy to adopt the techniques from Section 1.1 to show sharpness of the upper bound from Theorem 20 on the order-types for the iterations of reflection along elements of $\mathcal{O}$ necessary to attain the completeness of iterations of reflection for $\Pi_{2n+1}$-sentences.
In the present paper, we have been focusing on the iterations of full uniform reflection, however, when we want to obtain some true $\Pi_{n}$-sentence in principle it is natural to expect that we need to iterate just partial uniform $\Pi_{n}$ reflection $\mathsf{RFN}_{\Pi_{n}}$ instead of full uniform reflection $\mathsf{RFN}$. For example, the existence of such iterations could be easily proved by combining Feferman’s completeness theorem for full uniform reflection with Schmerl’s formula that allows us to transform short iterations of stronger reflection to longer iterations of weaker $\Pi_{n}$ reflection, while preserving the set of $\Pi_{n}$-consequences. We suspect that the methods used in this paper can be adopted to show that for some natural number constant $C$ and every $n$ we can replace $\mathsf{RFN}$ with $\mathsf{RFN}_{\Pi_{2n+1+C}}$ in Section 1.1. However, it is not clear to us whether in Section 1.1 full uniform reflection $\mathsf{RFN}$ can be replaced with just $\mathsf{RFN}_{\Pi_{2n+1}}$, while keeping the bound $\omega^{n}+1$ on the order-type.